# Holistic Explainable AI (H-XAI): Extending Transparency Beyond Developers in AI-Driven Decision Making

### v1

@@ 3.1 Rating-Driven Explanation (RDE)
This comparison grounds the explanation. A model that behaves closer to the biased baseline may raise fairness concerns; one that matches the random model may be unreliable. These comparisons are not just for evaluation: they help users understand why the model’s behavior is problematic and where. RDE outputs can be numeric scores or visual plots, depending on the user query. For example, Figure D.1 shows a case where RDE highlights age-based confounding in a credit classifier, while Figure D.3 shows how error rates change under missing input perturbations.
3.2 Traditional XAI Methods Used
While RDE provides explanations concerning the robustness of the models at a global level, it is not suitable for local feature attribution or communicating the robustness of the model at a global level. To address these needs, we integrate standard post-hoc explanation methods, which stakeholders may already be familiar with. The complete definitions of each can be found in Appendix C.
PDPs: Show the marginal effect of a feature on predictions by averaging over the dataset. Useful for understanding monotonicity or directionality, but unreliable when features are correlated. In Figure D.1, we use PDPs to examine the effect of credit amount across applicant groups.
SHAP: Estimates the marginal contribution of each feature to a prediction, using additive Shapley values. For time-series models, we use TreeSHAP on surrogate models (Raykar et al. 2023) to reduce computational cost. This makes the explanations more scalable, even for long input sequences. In our case study, we evaluated the surrogate on a small subset (10% of data, 60 samples) and found that its MASE and SMAPE differed from MOMENT’s by only 0.66 and 0.18%, respectively (see Table D.1 in the appendix).
Counterfactuals: Identify minimal input changes needed to flip a prediction (Wachter, Mittelstadt, and Russell 2017). This is useful for questions like, “What should this applicant change to get approved?”
3.3 Holistic-eXplainable AI (H-XAI)
@@
Figure 3: (a) Causal diagram for binary classification on the German Credit dataset. Protected attributes include Age, Personal Status, Gender, or their combinations, depending on query. (b) Causal diagram for time-series forecasting (stock prices), adapted from (Lakkaraju et al. 2024).
4 Case Studies
Figure 4: Hypothesis: Does decreasing the credit amount by 50% improve the chances of loan approval? Conclusion: On average, decreasing the requested credit amount by 50% increases the risk by 0.06, meaning applicants will be more likely to be classified as good risk (denoted by class 1). This suggests that asking for a smaller loan may improve the likelihood of loan approval only marginally.
@@ 4 Case Studies
Table 1: Mapping explanation methods to stakeholder questions, limitations, and when to combine them. Combinations show scenarios from our case studies (Section 4). Others are possible depending on task and data context.
@@
In this section, we present two case studies to demonstrate how H-XAI can address the explanation needs of different stakeholders (introduced in Section 2.1). These scenarios are structured around a taxonomy of questions that align with the XAI Question Bank (Liao et al. 2021). The mapping of our scenario questions to the categories of the XAI Question Bank is provided in the appendix (Table D.2). We show six scenarios: three for each stakeholder type in a given task. Three are included here, and the rest are in Appendix D.
4.1 Case Study-1: Binary Classification
@@
In the first case study, we choose binary credit risk classification as it is a well-studied, high-stakes domain where fairness, explanation, and model behavior have real-world consequences (Demajo, Vella, and Dingli 2020; Kłosok, Chlebus et al. 2020; Melsom, Vennerød, and de Lange 2022).
@@
A. Data. German Credit Dataset (Frank 2010) has been widely used in prior work on algorithmic fairness and explainability (Zemel et al. 2013; Martins et al. 2023; Verma and Rubin 2018), making it suitable for examining how different stakeholder questions surface in a binary classification task. It contains 1,000 rows with 20 applicant features, including demographic information (e.g., age, personal status and sex), financial details (e.g., credit amount, duration, checking account status), and other applicant-related attributes which are used to assess loan risk. We selectively analyze subsets of these features depending on the specific explanation or causal question being addressed. The primary outcome variable is binary: $Risk\in\{0,1\}$, where $Risk=1$ indicates a good credit risk (leading to loan approval), and $Risk=0$ indicates a bad credit risk (loan rejection).
@@
B. Causal Setup. For generating RDE, we use the causal diagram shown in Figure 3(a). Specifically, we focus on a smaller set of variables: we treat Credit Amount as the treatment, Age and Personal Status and Sex as protected attributes following (Zemel et al. 2013), and Risk as the outcome. In contrast, when applying XAI methods like SHAP, we do use the full feature set, since those methods are not bound by the same identifiability constraints and are designed to give feature attribution-based explanation rather than estimate causal effects.
@@
C. AI Models Considered. We use two models: a linear, interpretable model (Logistic Regression), and a non-linear ensemble model (Random Forest). They are two of the most popular methods used for credit scoring (Addy et al. 2024). Figure D.2 in the appendix highlights an accuracy versus confounding bias trade-off in Scenario-3. While Random Forest outperforms Logistic Regression in accuracy, it also exhibits higher confounding bias. This reflects the common, but not universal, trade-off between accuracy and interpretability (Atrey et al. 2025).
@@
We present three scenarios in which H-XAI (described in Section 3) is applied to address questions posed by different types of stakeholders: an individual applicant (Figure 5), a regulatory official (Figure D.1 in the appendix), and a data scientist within an operational organization (Figure D.2 in the appendix).
Figure 5: Stakeholder: Jack (applicant; individual); Scenario-1: Bank XYZ uses a Random Forest model to classify applicants as good or bad risk based on features like credit amount, age, personal status, gender, and more. Jack was classified as a bad risk and wants to understand why, and what changes he could make to be considered a good risk and get his loan approved.
@@
We also provide an example of a hypothesis-driven explanation for credit risk classification in Figure 4. The causal setup described for RDE in Section 3.1 also supports hypothesis-driven explanations, allowing users to test claims about model behavior rather than receive fixed outputs, as illustrated in Figure 4. In the appendix, we show how doubling age (Figure E.4) or halving loan duration (Figure E.5) affects predicted credit risk, demonstrating how hypothesis-driven explanations can serve as a standalone method for validating assumptions about model behavior.
@@
Figure 6: Stakeholder: Suresh (Regulatory official; Regulatory body). Scenario-4: Suresh uses a stock forecasting tool developed by a third-party company to guide her investment decisions. He notices that the tool performs poorly on certain days and suspects it may be less reliable for some companies. To decide whether to continue using it, he wants to understand how robust the tool is.
@@
Figure 7: Stakeholder: Jane (Retail Investor; Individual). Scenario-5: Jane uses a stock forecasting tool developed by a third-party company to guide her investment decisions. She notices that the tool performs poorly on certain days and suspects it may be less reliable for some companies. To decide whether to continue using it, she wants to understand how robust the tool is.
4.2 Case Study-2: Time-Series Forecasting
@@
A. Data. We collected daily stock prices from Yahoo! Finance for six companies across different industries: Meta (META) and Google (GOO) in social technology, Pfizer (PFE) and Merck (MRK) in pharmaceuticals, and Wells Fargo (WFC) and Citigroup (C) in financial services, as closing stock prices is the most commonly used data source for time-series forecasting (Lin and Marques 2024; Lakkaraju et al. 2024). The dataset covers from March 28, 2023, to April 22, 2024; data up to March 22, 2024, was used to forecast the following month.
B. Causal Setup. To generate RDE for time-series forecasting, we adopt the preprocessing approach from (Lakkaraju et al. 2024), where the difference between the 20 predicted stock prices (based on the previous 80 time steps) and their corresponding ground truth values is used to compute a sequence of residuals. While other metrics like SMAPE or MASE are available, we use residuals because they give the raw prediction error at each point, rather than at an aggregate level. To capture worst-case performance, we take the maximum residual within each prediction window as the outcome of interest. For simplicity, we refer to this as the residual throughout the paper. While predicted stock prices may vary across companies, our focus is on whether residuals, i.e., model errors, also vary systematically, which may indicate unreliability. This residual serves as the basis for analyzing how protected attributes or perturbations affect model behavior. Figure 3(b) shows the causal diagram used for this analysis. We treat Perturbation (e.g., introducing missing values or zero-imputation, as in Figure D.3) as the treatment, Company as the sensitive attribute, and Residual as the outcome. This causal setup is used exclusively for explanation questions that require RDE.
C. AI Models Considered. Our goal is not to benchmark forecasting performance but to analyze and explain model behavior. So rather than selecting the most competitive models, we chose two representative ones from the models used in (Lakkaraju et al. 2024): MOMENT, an open-source, time-series-specific foundation model, and Gemini, a general-purpose foundation model, both of which are easily accessible. We also use ARIMA, a statistical time-series forecasting model, as an additional baseline.
We present three scenarios in H-XAI and apply them to address questions posed by different types of stakeholders: a regulatory official (Figure 6), a retail investor (Figure 7), and a data scientist within an operational organization (Figure D.3 in the appendix).
4.3 When to Use: Mapping Questions to Methods
@@
5 Discussion and Conclusion
While our case studies focus on credit risk classification and stock price forecasting, H-XAI is task-agnostic and extends beyond these domains. It emphasizes selecting explanation techniques based on stakeholder roles and questions, rather than defaulting to a single method. In classification tasks, RDE may estimate the causal impact of applicant attributes on model decisions, while in forecasting, it can quantify how input perturbations affect performance. Traditional XAI methods complement RDE by providing local, instance-level explanations where causal analysis may not apply.
As shown in Figure 2, RDE is not just a metric computation step, but a structured reasoning process: constructing baselines, defining a causal model, and selecting appropriate causality-based metrics to answer the question. This structure enables H-XAI to adapt to different stakeholder needs, whether the goal is to understand robustness, quantify bias, or interpret individual predictions, making it suitable for critical domains like healthcare and finance. We also recognize that stakeholders may not always know how to define a causal model. To address this, RDE uses a generalized causal model by default, capturing common structural assumptions across tasks. Whether the analysis targets confounding bias (e.g., DIE %) or statistical bias (e.g., WRS) depends on how the question is framed, as demonstrated in Q1 of Figure 5 versus Q1 of Figure D.3. We do not treat interpretability as a fixed property of a method. Each explanation is shaped by the stakeholder’s question and the context.
A key takeaway from our work is that explanation is an iterative process, not a static output. No single method is sufficient on its own; different questions demand different tools. We show that by selecting and combining methods, whether RDE, counterfactuals, or feature attribution, we can shift the role of explanation from justifying predictions to supporting meaningful interaction, allowing stakeholders to test hypotheses and refine their understanding. While a promising step toward a unified approach, this work opens several avenues for future research, including expanded case studies, additional methods, and conducting user studies with diverse stakeholders.
@@ 5 Discussion and Conclusion
Appendix A Related Work
A.1 Explainability in AI
Explainable AI (XAI) is a set of processes and methods that are used to describe AI models, by characterizing their accuracy, transparency, and outcomes (Arrieta et al. 2020). AIX360 (Arya et al. 2019) introduces an open-source toolkit for explainability, featuring a range of SOTA techniques to guide users in selecting appropriate explanation methods based on their needs. While this toolkit is very valuable, the explanations do not allow users to interact with them. (Mueller et al. 2019) consider an effective explanation to be an interaction and not a property of statements. (Abdul et al. 2018) also acknowledges that a single explanation rarely results in a desired level of understanding. Explanation is an exploratory activity and never a ”one-off” (Hoffman et al. 2023a). (Chromik and Butz 2021) categorizes prior XAI work into different types of interaction based on the concepts introduced by (Hornbæk and Oulasvirta 2017). The paper highlights that prior research has been limited in facilitating natural, iterative conversations about AI behavior, an essential aspect for achieving transparency and scrutability in XAI systems. Additionally, such exploratory dialogues allow users to explore how changes in input lead to changes in the AI prediction (Chromik and Butz 2021). Through H-XAI, we address a series of user questions, where each question builds upon the previous one. This iterative process culminates in suggesting potential steps that the user can take to achieve the desired model outcome.
Feature importance methods such as SHapley Additive exPlanations (SHAP) and counterfactual explanations are the most widely-used explanation methods. The former is sensitive to feature correlations and the latter only explains a single outcome instead of the entire model (Alfeo et al. 2023). Relying on People benefit more from having local and global explanations that are integrated, yet the majority of XAI methods have focused on the delivery of local explanations (Hoffman et al. 2023b). We employ correlation-based XAI methods to provide high-level explanations of model decisions, while for more in-depth user questions that require understanding the underlying causal structure, we transition to causality-based explanations to offer deeper insights into the model’s behavior.
A.2 Explainability in Time-Series Forecasting Models (TSFMs)
Traditional XAI methods for TSFMs rely on feature attribution and visualization methods. Model-agnostic methods such as SHAP are widely used (Mokhtari, Higdon, and Başar 2019; Schlegel et al. 2019; Zhang et al. 2022), along with partial dependence plots (Greenwell 2017) to estimate marginal contributions of input variables and forecast outputs. These methods have been adapted to consider time-dependent inputs, but they are often limited in capturing causality or robustness under input perturbation (Lakkaraju et al. 2025). In deep learning-based TSFMs, attention mechanisms have also been used to show influential time steps or variables in multivariate forecasting tasks (Abbasimehr and Paki 2022; Li et al. 2019; Shih, Sun, and Lee 2019). These were used to picture instance-level temporal interpretability. However, such mechanisms are typically embedded within the models and therefore may not be reliable indicators to measure causal influence. Though these methods have been useful for prediction interpretability, they still tend to disproportionately focus on either local or global explanations as opposed to both. SHAP often provides local insights into single predictions. PDPs, for instance, offer aggregated trends across a dataset. In TSFMs, it is often necessary to combine both perspectives, i.e., understanding how individual time points contribute to certain forecasts (local), and how these patterns also generalize across entire data (global). Existing approaches do not always consistently support both forms of explanation in an integrated or interactive manner, especially under stakeholder-specific constraints. In recent years, there has been a growing shift in the literature toward methods that not only explain model predictions but also empower stakeholders at different levels to interrogate model behavior through hypothesis testing and counterfactual reasoning (Çelik, İcan, and Bulut 2023; Tasin et al. 2023). Motivated by this trend, our work supports stakeholder-aligned explanation for TSFMs (Section 4.2) across two key dimensions: accuracy and robustnes, which includes performance under varying input conditions as well as bias.
Appendix B RDE Metrics
Weighted Rejection Score (WRS): WRS quantifies statistical bias across a protected attribute $z_{i}\in Z$ by comparing outcome distributions $(O\mid z_{i})$ across all $\binom{m}{2}$ group pairs using Student’s t-test (Student 1908). For each pair, the t-value $t_{z_{i}}$ is compared to the critical value $t_{\text{crit}}$ for a given confidence interval (CI). A rejection indicates a significant difference in outcomes. WRS is computed as a weighted sum of rejections across CIs, where $x_{i}$ is the number of rejections and $w_{i}$ is the assigned weight (e.g., 1, 0.8, 0.6 for 95%, 75%, and 60% CIs). Formally, $WRS=\sum_{i}w_{i}\cdot x_{i}$.
We compute WRS when the treatment is applied uniformly across all protected attribute groups, making DIE % inappropriate in such cases (as in Section 4.2).
Average Treatment Effect (ATE): Average Treatment Effect (ATE) (Rubin 1974; Splawa-Neyman, Dabrowska, and Speed 1990) measures the average difference in outcomes between the treated ($T=i$) and untreated ($T=0$) units, quantifying the impact of the treatment on the AI model’s outcome. Unlike local counterfactuals, ATE quantifies the average effect of a treatment on model predictions across the dataset, making it well-suited for answering stakeholder questions about overall effect of an attribute on the outcome. Formally, ATE=[|E[O|do(T=i)]E[O|do(T=0)]|]ATE=[|E[O|do(T=i)]-E[O|do(T=0)]|]
Deconfounded Impact Estimation (DIE %): DIE % measures how much confounders ($Z$) distorts the relationship between a treatment ($T$) and an outcome ($O$), leading to confounding bias. We use Propensity Score Matching (PSM) to calculate DIE % when dealing with categorical treatments, but extend it to continuous treatments using G-computation (Robins 1986). Instead of linear models, we use a SuperLearner (Van der Laan, Polley, and Hubbard 2007), an ensemble of non-linear models to estimate counterfactual outcomes. DIE % evaluates the difference between the ATE computed without adjustment ($ATE_{0}$) and the ATE after applying G-computation ($ATE_{g}$). Formally, $DIE\%=\left[||ATE_{o}|-|ATE_{g}||\right]\times 100$

### v2

@@ 3.1.3. RDE Workflow
This comparison grounds the explanation. A model that behaves closer to the biased baseline may raise fairness concerns; one that matches the random model may be unreliable. These comparisons are not just for evaluation: they help users understand why the model’s behavior is problematic and where. RDE outputs can be numeric scores or visual plots, depending on the user query. For example, Figure 8 shows a case where RDE highlights age-based confounding in a credit classifier, while Figure 12 shows how error rates change under missing input perturbations.
3.2. Traditional XAI Methods Used
While RDE provides explanations concerning the robustness of the models at a global level, it is not suitable for local feature attribution or communicating the robustness of the model at a global level. To address these needs, we integrate standard post-hoc explanation methods, which stakeholders may already be familiar with.
Partial Dependency Plots (PDPs): Show the marginal effect of a feature on predictions by averaging over the dataset. Useful for understanding monotonicity or directionality, but unreliable when features are correlated. In Figure 8, we use PDPs to examine the effect of credit amount across applicant groups.
SHapley Additive exPlanations (SHAP): Estimates the marginal contribution of each feature to a prediction, using additive Shapley values. For time-series models, we use TreeSHAP on surrogate models (Raykar et al., 2023) to reduce computational cost. This makes the explanations more scalable, even for long input sequences. In our case study, we evaluated the surrogate on a small subset (10% of data, 60 samples) and found that its MASE and SMAPE differed from MOMENT’s by only 0.66 and 0.18%, respectively (see Table A.1 in the appendix).
Counterfactuals: Identify minimal input changes needed to flip a prediction (Wachter et al., 2017). This is useful for questions like, “What should this applicant change to get approved?”
3.3. Holistic eXplainable AI (H-XAI)
@@
Figure 3. (a) Causal diagram for binary classification on the German Credit dataset. Protected attributes include Age, Personal Status, Gender, or their combinations, depending on query. (b) Causal diagram for time-series forecasting (stock prices), adapted from (Lakkaraju et al., 2024a).
4. Case Studies
Figure 4. Hypothesis: Does decreasing the credit amount by 50% improve the chances of loan approval? Conclusion: On average, decreasing the requested credit amount by 50% increases the risk by 0.06, meaning applicants will be more likely to be classified as good risk (denoted by class 1). This suggests that asking for a smaller loan may improve the likelihood of loan approval only marginally.
Figure 5. Hypothesis: Does doubling applicants’ age, reduce their chances of being approved for a loan? Conclusion: On average, doubling a person’s age led to a noticeable drop in the predicted creditworthiness (risk score decreased by 0.35). This suggests that the model tends to classify older applicants as higher risk, which could be a sign of age-related bias.
Figure 6. Hypothesis: Does decreasing the loan duration by half improve the chances of loan approval? Conclusion: Reducing the loan duration by half led to a slight increase in predicted creditworthiness (risk score increased by 0.097). This indicates that the model tends to view shorter loan durations more favorably, which may slightly help an applicant get approved.
Figure 7. Stakeholder: Jack (applicant; individual); Scenario-1: Bank XYZ uses a Random Forest model to classify applicants as good or bad risk based on features like credit amount, age, personal status, gender, and more. Jack was classified as a bad risk and wants to understand why, and what changes he could make to be considered a good risk and get his loan approved.
Figure 8. Stakeholder: Maya (Regulatory official; Regulatory body). Scenario-2: An applicant reports that the AI model used by Bank XYZ is biased with respect to personal status and gender, and suspects their loan was rejected for this reason. Maya, a regulator from the ABC Bureau, begins an investigation to determine whether this bias is present.
Figure 9. Stakeholder: James (Data Scientist; Operational Organization). Scenario-3: Given a set of candidate models, James sets out to evaluate which one performs best on the client’s credit risk data. He plans to assess both accuracy and fairness before selecting a model for deployment; Interestingly, the global SHAP summary (Q2) is closely aligned with the Jack’s SHAP values shown in Figure 7!
@@ 4. Case Studies
Table 1. Mapping explanation methods to stakeholder questions, limitations, and when to combine them. Combinations show scenarios from our case studies (Section 4). Others are possible depending on task and data context.
@@
In this section, we present two case studies to demonstrate how H-XAI can address the explanation needs of different stakeholders (introduced in Section 2.1). These scenarios are structured around a taxonomy of questions that align with the XAI Question Bank (Liao et al., 2021). The mapping of our scenario questions to the categories of the XAI Question Bank is provided in the appendix (Table A.2). We show six scenarios: three for each stakeholder type in a given task.
4.1. Case Study-1: Binary Classification
@@
In the first case study, we choose binary credit risk classification as it is a well-studied, critical domain where fairness, explanation, and model behavior have real-world consequences (Demajo et al., 2020; Kłosok et al., 2020; Melsom et al., 2022).
@@
A. Data. German Credit Dataset (Frank, 2010) has been widely used in prior work on algorithmic fairness and explainability (Zemel et al., 2013; Martins et al., 2023; Verma and Rubin, 2018), making it suitable for examining how different stakeholder questions surface in a binary classification task. It contains 1,000 rows with 20 applicant features, including demographic information (e.g., age, personal status and sex), financial details (e.g., credit amount, duration, checking account status), and other applicant-related attributes which are used to assess loan risk. We selectively analyze subsets of these features depending on the specific explanation or causal question being addressed. The primary outcome variable is binary: $Risk\in\{0,1\}$, where $Risk=1$ indicates a good credit risk (leading to loan approval), and $Risk=0$ indicates a bad credit risk (loan rejection).
@@
B. Causal Setup. For generating RDE, we use the causal diagram shown in Figure 3(a). Specifically, we focus on a smaller set of variables: we treat Credit Amount as the treatment, Age and Personal Status and Sex as protected attributes following (Zemel et al., 2013), and Risk as the outcome. In contrast, when applying XAI methods like SHAP, we do use the full feature set, since those methods are not bound by the same identifiability constraints and are designed to give feature attribution-based explanation rather than estimate causal effects.
@@
C. AI Models Considered. We use two models: a linear, interpretable model (Logistic Regression), and a non-linear ensemble model (Random Forest). They are two of the most popular methods used for credit scoring (Addy et al., 2024). Figure 9 highlights an accuracy versus confounding bias trade-off in Scenario-3. While Random Forest outperforms Logistic Regression in accuracy, it also exhibits higher confounding bias. This shows the common, but not universal, trade-off between accuracy and interpretability (Atrey et al., 2025).
@@
We present three scenarios in which H-XAI (described in Section 3) is applied to address questions posed by different types of stakeholders: an individual applicant (Figure 7), a regulatory official (Figure 8), and a data scientist within an operational organization (Figure 9).
@@
We also provide an example of a hypothesis-driven explanation for credit risk classification in Figures 4, 5, 6. We define hypothesis-driven explanations as a structured process in which users test specific assumptions or causal questions about model behavior, rather than passively receiving fixed outputs. In our formulation, such hypotheses are operationalized through causal models, allowing users to quantify effects and systematically examine how the model behaves. While this is not exactly what Miller envisioned, it aligns with the direction of his Evaluative AI framework (Miller, 2023). The causal setup described for RDE in Section 3.1 also supports hypothesis-driven explanations, allowing users to test claims about model behavior rather than receive fixed outputs, as illustrated in Figure 4. We show how doubling age (Figure 5) or halving loan duration (Figure 6) affects predicted credit risk, demonstrating how hypothesis-driven explanations can serve as a standalone method for validating assumptions about model behavior.
@@
Figure 10. Stakeholder: Suresh (Regulatory official; Regulatory body). Scenario-4: Suresh uses a stock forecasting tool developed by a third-party company to guide her investment decisions. He notices that the tool performs poorly on certain days and suspects it may be less reliable for some companies. To decide whether to continue using it, he wants to understand how robust the tool is.
@@
Figure 11. Stakeholder: Jane (Retail Investor; Individual). Scenario-5: Jane uses a stock forecasting tool developed by a third-party company to guide her investment decisions. She notices that the tool performs poorly on certain days and suspects it may be less reliable for some companies. To decide whether to continue using it, she wants to understand how robust the tool is.
Figure 12. Stakeholder: Amanda (Data scientist; Operational organization). Scenario-6: Amanda is evaluating several time-series forecasting models for deployment. She wants to compare their accuracy and bias to choose one that best fits her team’s priorities.
4.2. Case Study-2: Time-Series Forecasting
@@
A. Data. We collected daily stock prices from Yahoo! Finance for six companies across different industries: Meta (META) and Google (GOO) in social technology, Pfizer (PFE) and Merck (MRK) in pharmaceuticals, and Wells Fargo (WFC) and Citigroup (C) in financial services, as closing stock prices is the most commonly used data source for time-series forecasting (Lin and Marques, 2024; Lakkaraju et al., 2024a). The dataset covers from March 28, 2023, to April 22, 2024; data up to March 22, 2024, was used to forecast the following month.
B. Causal Setup. To generate RDE for time-series forecasting, we adopt the preprocessing approach from (Lakkaraju et al., 2024a), where the difference between the 20 predicted stock prices (based on the previous 80 time steps) and their corresponding ground truth values is used to compute a sequence of residuals. While other metrics like SMAPE or MASE are available, we use residuals because they give the raw prediction error at each point, rather than at an aggregate level. To capture worst-case performance, we take the maximum residual within each prediction window as the outcome of interest. For simplicity, we refer to this as the residual throughout the paper. While predicted stock prices may vary across companies, our focus is on whether residuals, i.e., model errors, also vary systematically, which may indicate unreliability. This residual serves as the basis for analyzing how protected attributes or perturbations affect model behavior. Figure 3(b) shows the causal diagram used for this analysis. We treat Perturbation (e.g., introducing missing values or zero-imputation, as in Figure 12) as the treatment, Company as the sensitive attribute, and Residual as the outcome. This causal setup is used exclusively for explanation questions that require RDE.
C. AI Models Considered. Our goal is not to benchmark forecasting performance but to analyze and explain model behavior. So rather than selecting the most competitive models, we chose two representative ones from the models used in (Lakkaraju et al., 2024a): MOMENT, an open-source, time-series-specific foundation model, and Gemini, a general-purpose foundation model, both of which are easily accessible. We also use ARIMA, a statistical time-series forecasting model, as an additional baseline.
We present three scenarios in H-XAI and apply them to address questions posed by different types of stakeholders: a regulatory official (Figure 10), a retail investor (Figure 11), and a data scientist within an operational organization (Figure 12).
4.3. When to Use: Mapping Questions to Methods
@@
5. Discussion and Conclusion
Although our case studies focus on credit risk classification and stock price forecasting, H-XAI is task-agnostic and extends beyond these domains. The framework centers on selecting explanation techniques based on stakeholder roles and the questions they bring, which is increasingly important as AI systems are deployed through web-based decision services and online platforms. In classification settings, RDE can estimate the causal impact of applicant attributes on model decisions; in forecasting settings, it can quantify how input perturbations affect a model’s stability. Traditional XAI methods complement RDE by providing instance-level explanations where causal analysis may not apply.
As shown in Figure 2, RDE is more than a collection of metrics: it provides a step-by-step reasoning process that involves constructing appropriate baselines, outlining a causal model, and choosing the corresponding estimands. This structure allows H-XAI to adapt to a variety of stakeholder needs, whether the goal is to investigate robustness, quantify bias, or interpret individual decisions. These use cases naturally arise in settings where users interact with AI predictions through online interfaces, financial dashboards, or automated decision portals. We recognize that stakeholders may not always know how to specify a causal model. To handle this, RDE adopts a generalized causal model by default, capturing structural relationships that appear across many tasks. Depending on how a question is framed, the analysis may target confounding bias (e.g., DIE%) or statistical bias (e.g., WRS), as illustrated in Q1 of Figure 7 versus Q1 of Figure 12. Throughout the framework, interpretability is not treated as an inherent property of a single method but as something that emerges from aligning the method with the stakeholder’s specific question and context.
A core insight from this work is that explanation is an iterative process rather than a one-off output. Different questions require different tools, and no individual method is adequate on its own. By combining RDE with counterfactual reasoning and feature-based explanations, we shift the role of explanation from post hoc justification to a form of structured interaction that allows stakeholders to test hypotheses and improve their understanding of model behavior. While this paper presents a first step toward a holistic approach, several directions remain open for future work, including additional case studies, exploring complementary explanation methods, and conducting user studies with diverse stakeholders who rely on AI-driven decisions in online and operational settings.
@@ 5. Discussion and Conclusion
Appendix A RDE Metrics